# A Construction of Infinite Families of Self-Orthogonal Quasi-Cyclic Codes Using Constituent Codes

### v3

@@ 3.1.1. Recursive Construction:
$R_{1,2,...,2t+s}(\mathcal{C}_{1})$ $=$ $(d(C_{11}^{\prime})-d(C_{11}^{{}^{\prime}\perp_{E}}))d(D_{1})+(d(C_{11}^{{}^{\prime}\perp_{E}})-d(C_{12}^{\prime}))d(D_{1,2})+\ldots+d(C_{1s})d(D_{1,2,...,2t+s})$
$\geq$ $d(C_{1s})=d_{1s}.$
By Theorem 2.9, this shows $d(\mathcal{C}_{1})\geq d_{GO}(\mathcal{C}_{1})=\min\{R_{2t+s}(\mathcal{C}_{1}),R_{2t+s-1,2t+s}(\mathcal{C}_{1}),\ldots,R_{1,2,...,2t+s}(\mathcal{C}_{1})\}\geq d_{1s}$.
Now assume that for $u\in\mathbb{N}$, $d(\mathcal{C}_{u})\geq m^{u-1}d_{1s}$. Let $\mathcal{C}_{u+1}$ be the $m^{u+1}\ell$-length QC code constructed from $\mathcal{C}_{u}$, with constituent-code decomposition (46) and whose last constituent code is $\mathcal{C}_{u}$. Additionally, the minimum distances of the constituent codes of $\mathcal{C}_{u+1}$ over $\dfrac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}$ must satisfy equation (47), namely
@@
At this point, we do not have information on how $d(\mathcal{C}_{u})$ compares with the minimum distances of the other constituent codes. As the minimum distances in any decomposition of a QC code should satisfy the descending chain in  (25) to compute the GO-bound, we have the following $2t+s$ cases:
@@
For the first 2t+s12t+s-1 cases, we notice that all terms in the $R_{I}(\mathcal{C}_{u+1})$ in (24) are positive, and in particular, the last terms of $R_{2t+s}(\mathcal{C}_{u+1}),$ $R_{2t+s-1,2t+s}(\mathcal{C}_{u+1}),...,R_{1,2,...,2t+s}(\mathcal{C}_{u+1})$ are multiples of $d(C_{u+1,s-1})$. As, $d(C_{u+1,s-1})\geq m^{u}d(C_{1,s-1})\geq m^{u}d(C_{1,s})$. Thus for these cases, $d(\mathcal{C}_{u})\geq d(C_{u+1,s-1})\geq m^{u}d_{1s}$.
@@
$R_{2t+s}(\mathcal{C}_{u+1})$ $=$ $d(\mathcal{C}_{u})d(D_{2t+s})\geq m^{u-1}d_{1s}m=m^{u}d_{1s},\mbox{ from induction hypothesis,}$
$R_{2t+s-1,2t+s}(\mathcal{C}_{u+1})$ $=$ $(d(C_{u+1,s-1})-d(\mathcal{C}_{u}))d(D_{2t+s-1})+d(\mathcal{C}_{u})d(D_{2t+s-1,2t+s})$
$=$ $d(C_{u+1,s-1})d(D_{2t+s-1})-d(\mathcal{C}_{u})\underbrace{(d(D_{2t+s-1})-d(D_{2t+s-1,2t+s}))}_{>0}$
$\geq$ $d(C_{u+1,s-1})d(D_{2t+s-1})-d(C_{u+1,s-1})(d(D_{2t+s-1})-d(D_{2t+s-1,2t+s}))$
$=$ $d(C_{u+1,s-1})d(D_{2t+s-1,2t+s})$
$\geq$ $m^{u}d_{1s},$
@@
$R_{1,2,...,2t+s}(\mathcal{C}_{u+1})$ $=$ $(d(C_{u1}^{\prime})-d(C_{u+1,1}^{{}^{\prime}\perp}))d(D_{1})+\ldots+(d(C_{u+1,s-1})-d(\mathcal{C}_{u}))d(D_{1,2,...,2t+s-1})$
$+d(\mathcal{C}_{u})d(D_{1,2,...,2t+s})$
$=$ $\underbrace{(d(C_{u+1,1}^{\prime})-d(C_{u+1,1}^{{}^{\prime}\perp}))d(D_{1})+\ldots}_{\geq 0}+d(C_{u+1,s-1})d(D_{1,2,...,2t+s-1})$
- $d(\mathcal{C}_{u})\underbrace{(d(D_{1,2,...,2t+s-1})-d(D_{1,2,...,2t+s}))}_{>0}$
$\geq$ $d(C_{u+1,s-1})d(D_{1,2,...,2t+s-1})-d(C_{u+1,s-1})(d(D_{1,2,...,2t+s-1})-d(D_{1,2,...,2t+s}))$
$=$ $d(C_{u+1,s-1})d(D_{1,2,...,2t+s})$
$\geq$ $m^{u}d_{1s}$
where $d(D_{1,2,...,2t+s-1})-d(D_{1,2,...,2t+s})>0$ from Equation (4.19) in [cem]. Therefore by Theorem 2.9,
$d(\mathcal{C}_{u+1})\geq d_{GO}(\mathcal{C}_{u+1})=\min\{R_{2t+s}(\mathcal{C}_{u+1}),R_{2t+s-1,2t+s}(\mathcal{C}_{u+1}),\ldots,R_{1,2,...,2t+s}(\mathcal{C}_{u+1})\}\geq m^{u}d_{1s},$
as desired. ∎
The following theorem provides a summary of our construction of a QC code in $\mathbb{F}_{q}^{m^{u}\ell}$ that is self-orthogonal with respect to the Euclidean inner product.
@@
Let
$R_{q,m}^{\ell}\cong\bigoplus_{j=1}^{t}\left(\dfrac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\dfrac{\mathbb{F}_{q}[x]}{\langle g^{*}_{j}\rangle}\right)^{\ell}\oplus\bigg(\bigoplus_{i=1}^{s}\dfrac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\bigg)^{\ell},$
where $(g_{j},g_{j}^{*})$ are pairs of reciprocal irreducible factors, for $1\leq j\leq t$, and $f_{i}$ are self-reciprocal irreducible factors, for $1\leq 1\leq s$, with $f_{s}=x-1$. Let $\mathcal{C}_{u}$ be the $m^{u}\ell$-length QC code whose constituent codes are in (40) and (41), where $C_{ui}=m^{u-1}C_{1i}$, for all $1\leq i\leq s-1$. Then, we have an infinite family $\{\mathcal{C}_{u}\}_{u\geq 1}$ of ESO $q$-ary $\ell$-QC codes with parameters
(48) $\left[m^{u}\ell,\ell\left(\frac{m^{u}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+u\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}\right)+k_{1s},\geq m^{u-1}d_{1s}\right].$
Proof.
The parameters and duality of $\mathcal{C}_{u}$ comes from the recursive construction, and Lemmas 3.13.2. ∎
The next result is a particular case of Theorem 3.3, namely, where we remove some restrictions on the constituent codes which allow us also to construct infinite families of not only ESO QC codes, but also ones that are ESD and EDC. From now on, we assume $m$ is not even, when $q$ is odd (See Remark 1). In these special cases, the unique self-reciprocal irreducible factor of $x^{m}-1$ is $x-1\in\mathbb{F}_{q}[x]$. Hence, we no longer need to use copies of the constituent codes of $\mathcal{C}_{u}$ to obtain the constituent codes of $\mathcal{C}_{u+1}$. Consequently, we have no dimensional restriction on these codes, which allows for them to be ESD/EDC.
@@
Let
$R_{q,m}^{\ell}\cong\left(\bigoplus_{i=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{i}\rangle}\ \oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{i}^{*}\rangle}\right)^{\ell}\right)\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{\ell},$
where $(g_{i},g_{i}^{*})$ are pairs of distinct reciprocal polynomials, for all $1\leq i\leq t$. Let $\mathcal{C}_{1}\subset\mathbb{F}_{q}^{m\ell}$ be the ESO (resp. EDC or ESD) $\ell$-QC code with constituent codes $C_{1i}\oplus C_{1i}^{\perp}\subset\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{i}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{i}^{*}\rangle}\right)^{\ell}$ and a ESO (resp. EDC or ESD) $\ell,k_{1,t+1},d_{1,t+1}]_{q}$ code $C_{1,t+1}\subset\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{\ell}$, where $d(C_{11})\geq d^{\perp}(C_{11})\geq...\geq d(C_{1t})\geq d^{\perp}(C_{1t})\geq d(C_{1,t+1}).$ Then there exists an infinite family $\{\mathcal{C}_{u}\}_{u\geq 1}$ of ESO (resp. EDC or ESD) $q$-ary QC codes with parameters
(49) $\left[m^{u}\ell,\frac{\ell}{2}\left(m^{u}-1\right)+k_{1,t+1},\geq m^{u-1}d_{1,t+1}\right].$
Proof.
@@
(50) $\sum_{j=1}^{t}\deg(g_{j})=\frac{1}{2}\left[(m-\left(\sum_{i=1}^{s-1}\deg(f_{i})\right)-1\right]\overset{s=1}{=}\frac{1}{2}(m-1),$
@@
The duality of all constituent codes may be taken under the Euclidean inner product (see Remark 1 - item (1)) and it is defined exclusively by the $q$-ary constituent code $C_{1,t+1}$. Indeed, if $C_{1,t+1}$ is ESO (resp. EDC or ESD), then $\mathcal{C}_{1}$ also is. By the recursive construction, for any $u\geq 1$. the duality of $\mathcal{C}_{1}$ defines the duality of $\mathcal{C}_{2}$, and so on. This avoids relying on copies of constituent codes, which would otherwise restrict the construction to only ESO codes because of the dimensional limitations of self-dual and dual-containing codes. ∎
Remark 3.
We note that if we want $x^{m}-1$ to have three distinct irreducible factors in $\mathbb{F}_{q}[x]$, of which x1x-1 is one, we need that $m$ is prime or the square of a prime, in addition to $m$ and $q$ being coprime. Otherwise, $m$ would have at least two distinct prime factors so that in the factorization
(51) $x^{m}-1=\prod_{d\mid m}\Phi_{d}(x),$
where $\Phi_{d}$ is the $d$-th cyclotomic polynomial, there would be at least four cyclotomic polynomial factors ($1$, $m$, and $m$’s distinct prime factors). This would result in at least four irreducible factors. If $m$ is prime,
(52) $x^{m}-1=\Phi_{1}(x)\Phi_{m}(x)=(x-1)\Phi_{m}(x),$
so we must consider the irreducible factors of $\Phi_{m}$. The polynomial $\Phi_{m}$ has $\frac{\phi(m)}{\operatorname{ord}_{m}(q)}=\frac{m-1}{\operatorname{ord}_{m}(q)}$ irreducible factors. Thus, if $x^{m}-1$ has three irreducible factors, we need $\operatorname{ord}_{m}(q)=\frac{m-1}{2}$, i.e. $q^{(m-1)/2}\equiv 1\ (\operatorname{mod}\ m)$. If $m$ is the square of a prime, say $p^{2}$ where $p$ is prime,
(53) $x^{m}-1=\Phi_{1}(x)\Phi_{p}(x)\Phi_{p^{2}}(x).$
We need that both $\Phi_{p}$ and $\Phi_{p^{2}}$ are both irreducible, i.e., $1=\frac{\phi(p)}{\operatorname{ord}_{p}(q)}=\frac{p-1}{\operatorname{ord}_{p}(q)}$ so that $\operatorname{ord}_{p}(q)=p-1$ and $1=\frac{\phi(p^{2})}{\operatorname{ord}_{p^{2}}(q)}=\frac{p(p-1)}{\operatorname{ord}_{p^{2}}(q)}$ so that ${\operatorname{ord}_{p^{2}}(q)}=p(p-1)$. The following tables provide some examples of $m$ and $q$ pairs where $x^{m}-1\in\mathbb{F}_{q}[x]$ decomposes into one pair of reciprocal polynomials and x1x-1. One caveat is that over $\mathbb{F}_{q^{2}}[x]$, i.e., when $q$ is the square of a prime, $\Phi_{p}$ is irreducible if and only if $p=2$ by Fermat’s little theorem, and so $x^{p^{2}}-1$ does not have three irreducible factors over $\mathbb{F}_{q^{2}}$.
@@
Table 1. Some values of $m$ and $q$ for which $x^{m}-1\in\mathbb{F}_{q}[x]$ decomposes into three irreducible factors over $\mathbb{F}_{q}$.
@@
Table 2. Some values of $m$ and $q$ for which $x^{m}-1\in\mathbb{F}_{q^{2}}[x]$ decomposes into three irreducible factors over $\mathbb{F}_{q^{2}}$.

### v4

@@ 3.3 A Lower Bound for the Minimum Distance of the Family of Codes $\{\mathcal{C}_{u}\}_{u\geq 1}$
$N_{1,2,...,2t+s}(\mathcal{C}_{u})$ $=d(C_{u,1})d(D_{1,2,...,2t+s})\geq d(C_{u,2t+s})\geq m^{u-1}d(C_{1,s})=m^{u-1}d_{1,s}.$
Thus, $d(\mathcal{C}_{u})\geq d_{\mathrm{J}}(\mathcal{C}_{u})=\min\{N_{2t+s}(\mathcal{C}_{u}),N_{2t+s-1,2t+s}(\mathcal{C}_{u}),\ldots,N_{1,2,\ldots,2t+s}(\mathcal{C}_{u})\}\geq m^{u-1}d_{1,s}$ as desired. $\qed$
Example 3.23.
For sake of clarity, consider the following table of the decompositions of $\mathcal{C}_{u}$ for $2\leq u\leq 4$, where the decomposition is found in (29) and the constituent codes satisfy (34).
$u$ Decomposition
22 $\mathcal{C}_{2}\cong C_{2,1}\oplus C_{2,2}\oplus\cdots\oplus C_{2,2t+s}$
33 $\mathcal{C}_{3}\cong C_{3,1}\oplus C_{3,2}\oplus\cdots\oplus C_{3,2t+s}$
44 $\mathcal{C}_{4}\cong C_{4,1}\oplus C_{4,2}\oplus\cdots\oplus C_{4,2t+s}$
Table 1: Decompositions of $C_{2},\mathcal{C}_{3},\mathcal{C}_{4}$ to obtain lower bound for $d(\mathcal{C}_{4})$
Traversing horizontally through the row, we have that for $I=\{r,r+1,\ldots,2t+s\}\subseteq\{1,2,\ldots,2t+s\}$, $N_{I}(\mathcal{C}_{4})\geq d(C_{4,r})\geq d(C_{4,2t+s})$ by the descending chain condition, and then by equation  (34), traversing vertically up the table along the last constituent codes yields $N_{I}(\mathcal{C}_{4})\geq m^{3}d(C_{1,s})=m^{3}d_{1,s}$ for all $I$. Hence, $d(\mathcal{C}_{4})\geq m^{3}d_{1,s}.$
To obtain our lower bound, we include the condition that for each $u\geq 2$, and d(C’_uj^⟂_E) ≥md(C’_u-1,j^⟂_E), the last HSO constituent codes $C_{ui}$ in (28) satisfy
$d(C_{ui})\geq md(C_{u-1,i}),$ (36)
for all $1\leq i\leq s-1$. Note that repeated application of (34) yields $d(C_{ui})\geq m^{u-1}d(C_{1i})$ for each $i$.
proof We again use the induction argument along with the associated cyclic codes per (2.2). We first note that $D_{2t+s}$ is the $m$-length repetition code, so $d(D_{2t+s})=m$. For $u=1$, equation (16) and the fact that all nonzero constituent codes of $\mathcal{C}_{1}$ satisfy (10) yields
$N_{2t+s}(\mathcal{C}_{1})$ $=$ $d(C_{1s})d(D_{2t+s})=md_{1s}\geq d_{1s},$
$N_{2t+s-1,2t+s}(\mathcal{C}_{1})$ $=$ $d(C_{1,s-1})d(D_{2t+s-1,2t+s})\geq d(C_{1,s-1})\geq d_{1s},$
$\vdots$
$N_{1,2,...,2t+s}(\mathcal{C}_{1})$ $=$ $d(C^{\prime}_{11})d(D_{1,2,...,2t+s})\geq d(C_{1s})=d_{1s}.$
By Theorem 4, this shows $d(\mathcal{C}_{1})\geq\min\{N_{2t+s}(\mathcal{C}_{1}),N_{2t+s-1,2t+s}(\mathcal{C}_{1}),\ldots,N_{1,2,...,2t+s}(\mathcal{C}_{1})\}=d_{1s}$.
Now assume that for $u\in\mathbb{N}$, $d(\mathcal{C}_{u})\geq m^{u-1}d_{1s}$. Let $\mathcal{C}_{u+1}$ be the $m^{u+1}\ell$-length QC code constructed from $\mathcal{C}_{u}$, with constituent-code decomposition (3.20) and whose last constituent code is $\mathcal{C}_{u}$. At this point, we do not have information on how $d(\mathcal{C}_{u})$ compares with the minimum distances of the other constituent codes. As the minimum distance in any decomposition of a QC code should satisfy the descending chain in (10) to compute the Jensen bound, we consider the following two cases:
  • Case 1: $d(\mathcal{C}_{u})$ is not the smallest minimum distance.
    In this case, we notice that all terms
    $N_{2t+s}(\mathcal{C}_{u+1}),N_{2t+s-1,2t+s}(\mathcal{C}_{u+1}),...,N_{1,2,...,2t+s}(\mathcal{C}_{u+1})$
    are greater than or equal $d(C_{u+1,s-1})$ by definition of $N_{I}$, for $I\subseteq\{1,2,\ldots,2t+s\}$, in Theorem 16 and by the descending chain (10) for the constituent codes of $\mathcal{C}_{u+1}$. By (35), $d(C_{u+1,s-1})\geq m^{u}d(C_{1,s-1})$, which yields $m^{u}d(C_{1,s-1})\geq m^{u}d(C_{1,s})$ by the descending chain condition. Hence, $d(\mathcal{C}_{u+1})\geq d(C_{u+1,s-1})\geq m^{u}d_{1s}$.
  • Case 2: $d(\mathcal{C}_{u})$ is the smallest minimum distance
    For this case, we have
    $N_{2t+s}(\mathcal{C}_{u+1})$ $=$ $d(\mathcal{C}_{u})d(D_{2t+s})\geq m^{u-1}d_{1s}m=m^{u}d_{1s},\mbox{ from induction hypothesis,}$
    $N_{2t+s-1,2t+s}(\mathcal{C}_{u+1})$ $=$ $d(C_{u+1,s-1})d(D_{2t+s-1,2t+s})\geq m^{u}d(\mathcal{C}_{1,s-1})\geq m^{u}d_{1,s},$
    $\vdots$
    $N_{1,2,...,2t+s}(\mathcal{C}_{u+1})$ $=$ $d(C_{u+1,1}^{\prime})d(D_{1,2,...,2t+s})\geq d(C_{u+1,s-1})\geq m^{u}d(C_{1,s-1})\geq m^{u}d_{1s}.$
    Therefore by Theorem 4,
    $d(\mathcal{C}_{u+1})\geq\min\{N_{2t+s}(\mathcal{C}_{u+1}),N_{2t+s-1,2t+s}(\mathcal{C}_{u+1}),\ldots,N_{1,2,...,2t+s}(\mathcal{C}_{u+1})\}\geq m^{u}d_{1s},$
Thus for each $u\geq 1$, $d(\mathcal{C}_{u})\geq m^{u-1}d_{1s}$, as desired. dditionally, the minimum distances of the constituent codes of $\mathcal{C}_{u+1}$ over $\dfrac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}$ must satisfy equation (34), namely
@@ 44 $\mathcal{C}_{4}\cong C_{4,1}\oplus C_{4,2}\oplus\cdots\oplus C_{4,2t+s}$
At this point, we do not have information on how $d(\mathcal{C}_{u})$ compares with the minimum distances of the other constituent codes.
Suggestion: Using the Jensen Bound, the only time we really have to be careful is when $\mathcal{C}_{u}$ is the last constituent code. Otherwise, the definition of the $N_{I}$ and the descending chain of minimum distances causes the lower bound of $d_{J}(\mathcal{C}_{u})$ to be straightforwardly true. Hence, should we just consider two major case (Case 1: $\mathcal{C}_{u}$ is not the last constituent code, and Case 2: $\mathcal{C}_{u}$ is the last. Separate into these cases means we do not need to explciitly write out all $2t+s$ cases like below. Thoughts? G.: It makes sense!
Table showing the cases for where $C_{u}$ lives in $C_{u+1}$ according to the minimum distance descending chain condition:
$\mathcal{C}_{1}$
$\mathcal{C}_{2}$
$\mathcal{C}_{3}$
$\vdots$
$\mathcal{C}_{u+1}$
$\qed$
s the minimum distances in any decomposition of a QC code should satisfy the descending chain in  (10) to compute the GO-bound, we have the following $2t+s$ cases:
@@
For the first 2t+s12t+s-1 cases, we notice that all terms $N_{2t+s}(\mathcal{C}_{u+1}),$ $N_{2t+s-1,2t+s}(\mathcal{C}_{u+1}),...,N_{1,2,...,2t+s}(\mathcal{C}_{u+1})$ are greater than or equal $d(C_{u+1,s-1})$ by definition of $N_{I}$ for $I\subseteq\{1,2,\ldots,2t+s\}$ in Theorem 16 and by the descending chain of minimum distances for the constituents of $\mathcal{C}_{u+1}$. As $d(C_{u+1,s-1})\geq m^{u}d(C_{1,s-1})\geq m^{u}d(C_{1,s})$, then $d(\mathcal{C}_{u})\geq d(C_{u+1,s-1})\geq m^{u}d_{1s}$.
@@
$N_{2t+s}(\mathcal{C}_{u+1})$ $=$ $d(\mathcal{C}_{u})d(D_{2t+s})\geq m^{u-1}d_{1s}m=m^{u}d_{1s},\mbox{ from induction hypothesis,}$
$N_{2t+s-1,2t+s}(\mathcal{C}_{u+1})$ $=$ $d(C_{u+1,s-1})d(D_{2t+s-1,2t+s})\geq m^{u}d(\mathcal{C}_{1,s-1})\geq m^{u}d_{1,s},$
@@
$N_{1,2,...,2t+s}(\mathcal{C}_{u+1})$ $=$ $d(C_{u+1,1}^{\prime})d(D_{1,2,...,2t+s})\geq d(C_{u+1,s-1})\geq m^{u}d(C_{1,s-1})\geq m^{u}d(C_{1,s}).$
Therefore by Theorem 4,
$d(\mathcal{C}_{u+1})\geq\min\{N_{2t+s}(\mathcal{C}_{u+1}),N_{2t+s-1,2t+s}(\mathcal{C}_{u+1}),\ldots,N_{1,2,...,2t+s}(\mathcal{C}_{u+1})\}\geq m^{u}d_{1s},$
as desired.
The following theorem provides a summary of our construction of a family of ESO QC codes in $\mathbb{F}_{q}^{m^{u}\ell}$.
@@
Let $R_{q,m}^{\ell}$ be as in (25). Let $\mathcal{C}_{u}$ be the $m^{u}\ell$-length QC code whose constituent codes are in (27) and (28), where $C_{u,i}=m^{u-1}C_{1,i}$, for all $1\leq i\leq s-1$, and that their minimum distances satisfy (10). Then $\{\mathcal{C}_{u}\}_{u\geq 1}$ is an infinite family of ESO $q$-ary QC codes with parameters
$\left[m^{u}\ell,\ell\left(\frac{m^{u}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+u\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1,i}\right)+k_{,1s},\geq m^{u-1}d_{1,s}\right]_{q}.$ (37)
Proof 3.25.
The parameters and duality of $\mathcal{C}_{u}$ comes from the recursive construction, and Lemmas 3.19 and 3.21. $\qed$
The next result is a particular case of Theorem 3.24 which allow us also to construct infinite families of not only ESO (or EDC) QC codes, but also QC codes that are ESD. From now on, we assume $m$ is odd when $q$ is odd (see Remark 2.9) so that the unique self-reciprocal irreducible factor of $x^{m}-1$ is $x-1\in\mathbb{F}_{q}[x]$. Hence in these special cases, we no longer need to use copies of the constituent codes of $\mathcal{C}_{u}$ to obtain the constituent codes of $\mathcal{C}_{u+1}$. Consequently, as we no longer have the dimensional restriction of $\dim(C_{u,i})=\dim(C_{1,i})=k_{1,i}$ for each $1\leq i\leq s-1$, we may allow for the $\mathcal{C}_{u}$ to ESD. Note that $\ell$ must be even in order for $\mathcal{C}_{u}$ to be ESD.
@@
Let $R_{q,m}^{\ell}$ be as in (25) where the only self-reciprocal factor of $x^{m}-1$ is x1x-1. Let $\mathcal{C}_{1}\subset\mathbb{F}_{q}^{m\ell}$ be the ESO (resp. EDC or ESD) $\ell$-QC code whose constituent codes are $C_{1,j}\oplus C_{1,j}^{\perp_{E}}\subset\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\right)^{\ell}$ for $1\leq j\leq t$, and an ESO (resp. EDC or ESD) code $C_{1,t+1}\subset\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{\ell}$ with parameters $\ell,k_{1,t+1},d_{1,t+1}]_{q}$. Suppose that $d(C_{1,1})\geq d(C_{1,1}^{\perp_{E}})\geq\cdots\geq d(C_{1,t})\geq d(C_{1,t}^{\perp_{E}})\geq d(C_{1,t+1}).$ Then there exists an infinite family $\{\mathcal{C}_{u}\}_{u\geq 1}$ of ESO (resp. EDC or ESD) $q$-ary QC codes with parameters
$\left[m^{u}\ell,\ell\left(\frac{m^{u}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+k_{1,t+1},\geq m^{u-1}d_{1,t+1}\right]_{q}.$ (38)
Proof 3.27.
@@
$\sum_{j=1}^{t}\deg(g_{j})=\frac{1}{2}\left[(m-\left(\sum_{i=1}^{s-1}\deg(f_{i})\right)-1\right]=\frac{1}{2}(m-1),$ (39)
@@
The duality of all constituent codes may be taken under the Euclidean inner product (see Remark 2.9 - item (i)) and it is defined exclusively by the $q$-ary constituent code $C_{1,t+1}$. Indeed, if $C_{1,t+1}$ is ESO (resp. EDC or ESD), then $\mathcal{C}_{1}$ also is. By the recursive construction, for any $u\geq 1$, the duality of $\mathcal{C}_{1}$ defines the duality of $\mathcal{C}_{2}$, and so on. This avoids relying on copies of constituent codes, which would otherwise restrict the construction to only ESO (or EDC) codes because of the dimensional limitations of self-dual and dual-containing codes. $\qed$
We note that if we want $x^{m}-1$ to have three distinct irreducible factors in $\mathbb{F}_{q}[x]$, of which x1x-1 is one, we need that $m$ is prime or the square of a prime. Otherwise, $m$ would have at least two distinct prime factors so that in the factorization
$x^{m}-1=\prod_{d\mid m}\Phi_{d}(x),$ (40)
where $\Phi_{d}(x)$ denotes the $d$-th cyclotomic polynomial, there would be at least four cyclotomic polynomial factors ($1$, $m$, and the distinct prime factors of $m$). This would result in at least four irreducible factors.
If $m$ is prime,
$x^{m}-1=\Phi_{1}(x)\Phi_{m}(x)=(x-1)\Phi_{m}(x),$ (41)
so we must consider the irreducible factors of $\Phi_{m}$. The polynomial $\Phi_{m}$ has $\frac{\phi(m)}{\operatorname{ord}_{m}(q)}=\frac{m-1}{\operatorname{ord}_{m}(q)}$ irreducible factors. Thus, if $x^{m}-1$ has three irreducible factors, we need $\operatorname{ord}_{m}(q)=\frac{m-1}{2}$, i.e. $q^{(m-1)/2}\equiv 1\ (\operatorname{mod}\ m)$.
If $m$ is the square of a prime, say $p^{2}$ where $p$ is prime,
$x^{m}-1=\Phi_{1}(x)\Phi_{p}(x)\Phi_{p^{2}}(x).$ (42)
We need that both $\Phi_{p}$ and $\Phi_{p^{2}}$ are both irreducible, i.e., $1=\frac{\phi(p)}{\operatorname{ord}_{p}(q)}=\frac{p-1}{\operatorname{ord}_{p}(q)}$ so that $\operatorname{ord}_{p}(q)=p-1$ and $1=\frac{\phi(p^{2})}{\operatorname{ord}_{p^{2}}(q)}=\frac{p(p-1)}{\operatorname{ord}_{p^{2}}(q)}$ so that ${\operatorname{ord}_{p^{2}}(q)}=p(p-1)$. Tables 2 and 3 provide some examples of $m$ and $q$ pairs where $x^{m}-1\in\mathbb{F}_{q}[x]$ decomposes into one pair of reciprocal polynomials and x1x-1. One caveat is that over $\mathbb{F}_{q^{2}}[x]$, i.e., when $q$ is the square of a prime, $\Phi_{p}$ is irreducible if and only if $p=2$ by Fermat’s little theorem, and so $x^{p^{2}}-1$ does not have three irreducible factors over $\mathbb{F}_{q^{2}}$.
@@
Table 2: Some values of $m$ and $q$ for which $x^{m}-1\in\mathbb{F}_{q}[x]$ decomposes into three irreducible factors over $\mathbb{F}_{q}$.
@@
Table 3: Some values of $m$ and $q^{2}$ for which $x^{m}-1\in\mathbb{F}_{q^{2}}[x]$ decomposes into three irreducible factors over $\mathbb{F}_{q^{2}}$.